Over the weekend, a total of 8.5 containers were removed from the aft section of Rena. This includes 4.5 containers from hold 6 and 4 containers from hold 7.
The Smit Borneo is secured on the port quarter of Rena by wire ropes and positioned to work the aft holds and sea floor in the immediate vicinity.
The No.7 port side hatch cover was removed and wet stored at location.
53 big bags containing ‘lamb’ were discharged to the ST-60 from formerly refrigerated (reefer) containers.
One reefer unit was cut into two pieces and removed to the ST-60.
One container carrying ferro-silicon has been ‘wet stored’ in 9m of water near Rena after higher than normal gas levels were detected. Plans are in place for its removal.
Electronic sensor readings on the fore & aft section of Rena indicate no significant changes.
Svitzer plan to continue today with removal of cargo from the reefer units and cut them as required for disposal ashore.
Dive operations will continue (conditions permitting) in the aft section of the wreck, initially working in hold No.7 and then working towards shifting into No.5 starboard side, when possible.
The weather forecast for today is for Easterly winds of around 15 knots, changing later in the day back to the North and easing to light variable winds for Tuesday morning, rising to 16 knots on Tuesday afternoon and later increasing to 20-25 knot Northerlies Wednesday.
The sea state is expected to remain around 1.0m through until Tuesday. A strengthening northerly wind will elevate the sea state through Wednesday with around 2.0m significant wave heights currently expected by Wednesday afternoon / Thursday morning. Around 2.2m significant wave heights are expected on Thursday.
The Harbour Master reports there were 4 incursions into the exclusion zone over the weekend, which have been dealt with.
However, all vessel operators are reminded that there is still a 3NM water and air exclusion zone around and above the Rena for operational safety reasons.
The number of containers retrieved from the decks of Rena and brought to Port now stands at 544. The number retrieved from the sea and beaches remains 70.
A further 17 total have been located away from Rena, but are yet to be recovered.
Container debris has been removed from Bowentown.
Blue twine – enough to fill two skips – was collected from the ocean yesterday (Sunday). A helicopter was used to locate the floating twine which was then collected by fast response vessel and barge. The twine was found between Waihi Beach and Mayor Island.
Beach team members collected foam from the shore between Mount Maunganui main beach and Papamoa Domain yesterday. The foam was from refrigerated containers.
Braemar / Unimar barges were also used over the weekend to transport not only discharged containers, parts of containers, and twine but also 119 bulk bags of lamb that were moved from the Rena and sent straight to landfills.
Sonar sweeps continued over the weekend.
MNZ / Spill response
A clean up team of 16 will be working on Motiti Island this week.
Planning is continuing for operations on the East Coast and Matakana Island.
The last oil-affected birds were successfully released back into the wild on Friday. Monitoring over the weekend detected no problems with these birds. However, two un-oiled birds that appear to have died of natural causes were found.
The spill response team remains ready to scale up quickly should more oil be released from Rena.
The total waste collected remains at 1,044m3.
Source: Maritime NZ, February 20, 2012